OpenStudy (anonymous):

Hey! I need help with 3 questions! Please&Thank You! 1. If the hydrogen ion concentration of a solution is 10^-10 M is the solution acidic, alkaline, or nuetral? -acidic -alkaline -nuetral -cannot be determined 2. What is pH? -the negative logarithm of the hydrogen ion concentration -the positive logarithm of the hydrogen ion concentration -the negitive logarithm of the hydroxide ion concentration -the positive logarithm of the hydroxide ion concentration 3.Which of these solutions is the most basic? -[H+]=1*10^-2 M -[OH-]=1*10^-4 M -[H+]=1*10^-11 M -[OH-]= 1*10^-13 M

OpenStudy (aaronq):

Note: alkaline = basic Place on a continuum, this is what the pH scale looks like pH 1 pH=7 pH =14 acidic ->>>>>>>>>>>> neutral << <<<<<<<<<- basic from the equation \(pH=-log[H^+]\) you can verify that at pH=1; \([H^+]=0.1~M\) pH=7; \([H^+]=1.0*10^{-7}~M\) pH=14; \([H^+]=1.0*10^{-14}~M\)

OpenStudy (aaronq):

Any concentration above pH 7 is basic while below is acidic